# Foundation Model-Aided Hierarchical Control for Robust RIS-Assisted Near-Field Communications

### v1

@@ I INTRODUCTION
The evolution toward sixth-generation (6G) wireless systems will be driven by the demand for unprecedented data rates, ultra-low latency, and massive connectivity [saad2019vision]. A key enabler is the deployment of extremely large aperture arrays (ELAAs) operating in millimeter-wave (mmWave) and terahertz (THz) bands, which offer the massive spatial multiplexing gain required to meet 6G targets [rappaport2019wireless]. In high-frequency bands, small wavelengths enable practical deployment of ELAAs with hundreds of antenna elements. However, the combination of large apertures and high frequencies requires a paradigm shift from the far-field planar wave assumptions to the near-field communication (NFC) regime [dai2022near]. In NFC, signals propagate as spherical waves, making channel characteristics dependent on both angle and distance to the receiver [an2023codebook]. Given that the NFC in 6G systems can extend beyond 100 meters [bjornson2022massive], strictly angular beamsteering is insufficient. Instead, beamfocusing is required to precisely concentrate signal energy at specific spatial locations.
Realizing the potential of NFC is impeded by two primary physical constraints: severe propagation loss and spherical wavefront complexity. The mmWave and THz bands are susceptible to severe path loss and are easily blocked by physical objects [giordani2019survey]. Moreover, spherical wavefronts introduce significant nonlinearity into channel state information (CSI) estimation, requiring models that account for spatial variations across the large array aperture [cui2023near, he2023near]. This makes conventional CSI acquisition computationally prohibitive due to high pilot overhead [heath2016overview], yet accurate high-dimensional CSI is a prerequisite for effective beamfocusing [ma2021sparse].
@@ I-A1 NFC Channel Modeling and Estimation
The transition to ELAAs has necessitated a re-evaluation of channel models. Unlike far-field systems that assume planar wavefronts, near-field channels depend on both angle and distance, significantly increasing estimation complexity [tang2023nearfield]. Recent works have explored phenomena such as beam squint in wideband systems [zhang2023beamsquint]. However, accurately estimating these high-dimensional channels remains a bottleneck. While deep learning approaches have been suggested [wang2024transformer, elbir2020deep], traditional CNN-based estimators struggle with the long-range spatial dependencies inherent to large aperture arrays. Furthermore, few existing works explicitly leverage the geometric priors of near-field propagation to reduce computational complexity in transformer architectures, a gap our work addresses.
@@ I-A2 Blockage Prediction
Beyond channel estimation, link blockage poses a critical challenge. Traditional approaches such as relay switching are reactive. Recent research has shifted toward proactive strategies using multi-modal sensors [alkhateeb2018machine, charan2021vision]. Methods fusing multi-modal data have shown promise in building environment-aware networks [ghassemi2024multi]. Vision transformers (ViTs) demonstrate exceptional performance in blockage prediction by processing sequential image data [wang2022vision, ghassemi2025generative]. ViTs identify motion patterns of users and objects that indicate future blockages, offering superior generalization compared to classical signal processing methods [sanchez2022millimeter]. By integrating predictive capabilities, a network can preemptively reconfigure resources to maintain connectivity.
@@ I-A3 Hierarchical Control in Wireless Networks
Reinforcement learning (RL) has been widely applied to RIS optimization [ghassemi2025foundation]. However, RL agents suffer from the problem of dimensionality when controlling ELAAs. Hierarchical deep reinforcement learning (HDRL) offers a solution by decomposing timescales [huang2023hierarchical, liu2021deep]. Such decomposition is critical for RIS-assisted systems where beamforming must adapt to fast fading while phase shifts may be updated on a slower frame-based schedule [deng2022reconfigurable]. Our work extends this by decoupling the problem based on the information dynamics of CSI and blockage prediction results.
@@ I-B Motivations and Contributions
To address these challenges, we propose a dual-transformer (DT) hierarchical framework that integrates specialized transformer models within a HDRL architecture for proactive control in RIS-assisted NFC systems. Our approach coordinates proactive blockage prediction with real-time CSI estimation through a dual-timescale design that handles the inherent mismatch between slow blockage dynamics and fast channel fading. Recent advances in RIS-assisted systems have demonstrated the potential of intelligent reflecting surfaces for enhancing wireless coverage and capacity [liu2024ris], motivating our integration of foundation models with hierarchical control.
@@ II-A3 NFC Spherical Wave Channel Model
We adopt the spherical wave model which accounts for both phase variations and distance-dependent path loss across the array aperture [bjornson2022massive].
1) BS-to-UE Direct Channel: The distance from the $n$-th BS antenna to the $k$-th UE is $d_{k,n}^{\text{BD}}=\|\mathbf{u}_{k}-(\mathbf{u}_{\text{BS}}+\mathbf{p}_{n}^{\text{BS}})\|=\|\mathbf{u}_{k}-\mathbf{p}_{n}^{\text{BS}}\|$, where $\|\cdot\|$ denotes the Euclidean norm. The channel coefficient from the $n$-th BS antenna to UE $k$ is modeled as [bjornson2022massive]:
@@ II-B NFC Propagation Analysis
We analyze the phase discrepancy between near-field and far-field assumptions. The boundary between these regions is defined by the Rayleigh distance $Z_{R}$ [bjornson2022massive, dai2022near]:
@@
where $D$ is the maximum array aperture dimension. For ELAAs, $D$ is large relative to $\lambda$, causing $Z_{R}$ to extend significantly. For instance, with $N=1024$ antennas configured as $32\times 32$ at $d_{\text{BS}}=\lambda/2$ and $f_{c}=3.5$ GHz, the aperture $D\approx 1.88$ m yields $Z_{R}\approx 82$ m, placing typical urban deployment distances firmly in the near-field regime. This analysis applies generally to any ELAA (BS or RIS). For a user located at $\mathbf{u}_{k}$ with radial distance $r=\|\mathbf{u}_{k}\|$ and angle $\theta$ relative to the array boresight, let $q_{n}$ denote the generic scalar position of the $n$-th element along the array aperture. The phase of the signal under the planar wave assumption is $\phi_{\text{planar}}(n)=\frac{2\pi}{\lambda}q_{n}\sin(\theta)$. However, the exact spherical phase is determined by the Euclidean distance [bjornson2022massive]:
@@
A related challenge in wideband near-field systems is beam squint, where the beam direction varies across frequency subcarriers due to the wavelength-dependent phase term $e^{-j\frac{2\pi}{\lambda}d}$ [gao2023wideband]. Since $\lambda=c/f$ varies with frequency, each subcarrier in an OFDM system experiences a different phase shift, causing the focused beam to deviate spatially across the signal bandwidth. In the near-field regime, where distances $d_{k,n}$ vary significantly across the array aperture, this frequency-dependent beam deviation becomes severe, leading to signal distortion and power loss. Traditional far-field beamforming cannot compensate for this effect, as it assumes angle-only dependence without accounting for the distance-frequency coupling inherent to spherical wavefronts.
@@ IV-F SE Performance
Ablation Study (Theoretical): It is worth noting the contribution of the specific geometric features used in our CSI transformer. Standard far-field estimators rely solely on Angle of Arrival (AoA). In the NFC, AoA is distance-dependent. By explicitly embedding the distance $d_{k,n}$ as part of the input feature vector $\mathbf{x}_{k}$ (Eq. 15), our model linearizes the phase wavefront in the high-dimensional latent space. Without this distance embedding, the attention mechanism would fail to distinguish between users at the same angle but different depths, leading to significant “beam squint” (frequency-dependent beam deviation across subcarriers) losses [gao2023wideband]. The proposed architecture effectively acts as a non-linear phase compensator, validating its superiority over standard CNNs shown in Fig. 2.

### v2

@@ I INTRODUCTION
The evolution toward sixth-generation (6G) wireless systems will be driven by the demand for unprecedented data rates, ultra-low latency, and massive connectivity [saad2019vision]. A key enabler is the deployment of extremely large aperture arrays (ELAAs) operating in millimeter-wave (mmWave) and terahertz (THz) bands, which offer the massive spatial multiplexing gain required to meet 6G targets [rappaport2019wireless]. In high-frequency bands, small wavelengths enable practical deployment of ELAAs with hundreds of antenna elements. However, the combination of large apertures and high frequencies requires a paradigm shift from the far-field planar wave assumptions to the near-field communication (NFC) regime [cui2022near]. In NFC, signals propagate as spherical waves, making channel characteristics dependent on both angle and distance to the receiver [wei2022codebook]. Given that the NFC in 6G systems can extend beyond 100 meters [bjornson2020power], strictly angular beamsteering is insufficient. Instead, beamfocusing is required to precisely concentrate signal energy at specific spatial locations.
Realizing the potential of NFC is impeded by two primary physical constraints: severe propagation loss and spherical wavefront complexity. The mmWave and THz bands are susceptible to severe path loss and are easily blocked by physical objects [rangan2014millimeter]. Moreover, spherical wavefronts introduce significant nonlinearity into channel state information (CSI) estimation, requiring models that account for spatial variations across the large array aperture [cui2023near, lei2023channel]. This makes conventional CSI acquisition computationally prohibitive due to high pilot overhead [heath2016overview], yet accurate high-dimensional CSI is a prerequisite for effective beamfocusing [ma2020sparse].
@@ I-A1 NFC Channel Modeling and Estimation
The transition to ELAAs has necessitated a re-evaluation of channel models. Unlike far-field systems that assume planar wavefronts, near-field channels depend on both angle and distance, significantly increasing estimation complexity [cui2022channel]. Recent works have explored phenomena such as beam squint in wideband systems [zhang2022beam]. However, accurately estimating these high-dimensional channels remains a bottleneck. While deep learning approaches have been suggested [wang2022transformer, elbir2020deep], traditional CNN-based estimators struggle with the long-range spatial dependencies inherent to large aperture arrays. Furthermore, few existing works explicitly leverage the geometric priors of near-field propagation to reduce computational complexity in transformer architectures, a gap our work addresses.
@@ I-A2 Blockage Prediction
Beyond channel estimation, link blockage poses a critical challenge. Traditional approaches such as relay switching are reactive. Recent research has shifted toward proactive strategies using multi-modal sensors [alkhateeb2018machine, charan2021vision]. Methods fusing multi-modal data have shown promise in building environment-aware networks [ghassemi2024multi]. Vision transformers (ViTs) demonstrate exceptional performance in blockage prediction by processing sequential image data [charan2022computer, ghassemi2025generative]. ViTs identify motion patterns of users and objects that indicate future blockages, offering superior generalization compared to classical signal processing methods [yang2023hierarchical]. By integrating predictive capabilities, a network can preemptively reconfigure resources to maintain connectivity.
@@ I-A3 Hierarchical Control in Wireless Networks
Reinforcement learning (RL) has been widely applied to RIS optimization [ghassemi2025foundation]. However, RL agents suffer from the problem of dimensionality when controlling ELAAs. Hierarchical deep reinforcement learning (HDRL) offers a solution by decomposing timescales [zhou2022hierarchical, nguyen2022ris]. Such decomposition is critical for RIS-assisted systems where beamforming must adapt to fast fading while phase shifts may be updated on a slower frame-based schedule [zhi2022two]. Our work extends this by decoupling the problem based on the information dynamics of CSI and blockage prediction results.
@@ I-B Motivations and Contributions
To address these challenges, we propose a dual-transformer (DT) hierarchical framework that integrates specialized transformer models within a HDRL architecture for proactive control in RIS-assisted NFC systems. Our approach coordinates proactive blockage prediction with real-time CSI estimation through a dual-timescale design that handles the inherent mismatch between slow blockage dynamics and fast channel fading. Recent advances in RIS-assisted systems have demonstrated the potential of intelligent reflecting surfaces for enhancing wireless coverage and capacity [yuan2021reconfigurable], motivating our integration of foundation models with hierarchical control.
@@ II-A3 NFC Spherical Wave Channel Model
We adopt the spherical wave model which accounts for both phase variations and distance-dependent path loss across the array aperture [bjornson2020power].
1) BS-to-UE Direct Channel: The distance from the $n$-th BS antenna to the $k$-th UE is $d_{k,n}^{\text{BD}}=\|\mathbf{u}_{k}-(\mathbf{u}_{\text{BS}}+\mathbf{p}_{n}^{\text{BS}})\|=\|\mathbf{u}_{k}-\mathbf{p}_{n}^{\text{BS}}\|$, where $\|\cdot\|$ denotes the Euclidean norm. The channel coefficient from the $n$-th BS antenna to UE $k$ is modeled as [bjornson2020power]:
@@ II-B NFC Propagation Analysis
We analyze the phase discrepancy between near-field and far-field assumptions. The boundary between these regions is defined by the Rayleigh distance $Z_{R}$ [bjornson2020power, cui2022near]:
@@
where $D$ is the maximum array aperture dimension. For ELAAs, $D$ is large relative to $\lambda$, causing $Z_{R}$ to extend significantly. For instance, with $N=1024$ antennas configured as $32\times 32$ at $d_{\text{BS}}=\lambda/2$ and $f_{c}=3.5$ GHz, the aperture $D\approx 1.88$ m yields $Z_{R}\approx 82$ m, placing typical urban deployment distances firmly in the near-field regime. This analysis applies generally to any ELAA (BS or RIS). For a user located at $\mathbf{u}_{k}$ with radial distance $r=\|\mathbf{u}_{k}\|$ and angle $\theta$ relative to the array boresight, let $q_{n}$ denote the generic scalar position of the $n$-th element along the array aperture. The phase of the signal under the planar wave assumption is $\phi_{\text{planar}}(n)=\frac{2\pi}{\lambda}q_{n}\sin(\theta)$. However, the exact spherical phase is determined by the Euclidean distance [bjornson2020power]:
@@
A related challenge in wideband near-field systems is beam squint, where the beam direction varies across frequency subcarriers due to the wavelength-dependent phase term $e^{-j\frac{2\pi}{\lambda}d}$ [cui2024near]. Since $\lambda=c/f$ varies with frequency, each subcarrier in an OFDM system experiences a different phase shift, causing the focused beam to deviate spatially across the signal bandwidth. In the near-field regime, where distances $d_{k,n}$ vary significantly across the array aperture, this frequency-dependent beam deviation becomes severe, leading to signal distortion and power loss. Traditional far-field beamforming cannot compensate for this effect, as it assumes angle-only dependence without accounting for the distance-frequency coupling inherent to spherical wavefronts.
@@ IV-F SE Performance
Ablation Study (Theoretical): It is worth noting the contribution of the specific geometric features used in our CSI transformer. Standard far-field estimators rely solely on Angle of Arrival (AoA). In the NFC, AoA is distance-dependent. By explicitly embedding the distance $d_{k,n}$ as part of the input feature vector $\mathbf{x}_{k}$ (Eq. 15), our model linearizes the phase wavefront in the high-dimensional latent space. Without this distance embedding, the attention mechanism would fail to distinguish between users at the same angle but different depths, leading to significant “beam squint” (frequency-dependent beam deviation across subcarriers) losses [cui2024near]. The proposed architecture effectively acts as a non-linear phase compensator, validating its superiority over standard CNNs shown in Fig. 2.